# Corks, exotic 4-manifolds and genus functions

### v3

@@ 8.1. Submanifolds of $S^{4}(L)$
Before discussing genus functions, we show that various 4-manifolds are realized as codimension-00 submanifolds of the surgered 4-manifold $S^{4}(L)$ (see Definition 6.3) for some $S^{2}$-link $L$ in $S^{4}$.
@@
Let $L,L^{\prime}$ be $S^{2}$-links in $S^{4}$, and let $V,V^{\prime}$ be connected codimension-00 submanifolds of $S^{4}(L),S^{4}(L^{\prime})$ with non-empty boundaries, respectively. Then, the following hold. Note that $S^{4}(L)=S^{4}$ in the case where $L$ is an empty set.
$(1)$ The boundary sum $V\natural V^{\prime}$ admits an embedding into $S^{4}(L\sqcup L^{\prime})$, where $L\sqcup L^{\prime}$ denotes the disjoint union of $L$ and $L^{\prime}$ in $S^{4}\#S^{4}\cong S^{4}$.
$(2)$ Attach 1-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(3)$ Attach an algebraically canceling pair of $1$- and $2$-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(4)$ Attach $2$-handles to $V$ along a strongly slice link in $\partial V$ with 00-framings. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(5)$ Let $\ell$ be a link in $\partial(\natural_{n}S^{2}\times D^{2})$ that bounds a disjoint union of smoothly and properly embedded disks in $\natural_{n}S^{2}\times D^{2}$. Attach $2$-handles to $\natural_{n}S^{1}\times D^{3}$ along a link $\ell$ with the framings induced from the disks, where $\ell$ is regarded as a link in $\partial(\natural_{n}S^{1}\times D^{3})$ using an arbitrary diffeomorphism $\partial(\natural_{n}S^{2}\times D^{2})\to\partial(\natural_{n}S^{1}\times D^{3})$. Then the resulting $4$-manifold admits an embedding into $S^{4}$.
@@
Using the above lemma, we can generate numerous examples. Here we provide simple examples.
@@
Let $\Sigma_{g}$ denote a closed surface of genus $g$. Then $\Sigma_{g}\times D^{2}$ $(g\geq 0)$, $\natural_{n}S^{1}\times D^{3}$ $(n\geq 0)$, Mazur-type contractible 4-manifolds, their boundary sums, and their connected sums admit embeddings into $S^{4}$ and hence realized as codimension-00 submanifolds of $S^{4}$. Furthermore, these 4-manifolds with 2-handles attached along strongly slice links with 0-framings are realized as codimension-00 submanifolds of $S^{4}$.
@@
$(2)$ For any codimension-00 submanifold of $S^{4}$, its second homology group is torsion-free.
$(3)$ If the boundary of a codimension-00 submanifold $Z$ of $S^{4}(L)$ is a disjoint union of rational homology $3$-spheres, then $H_{2}(Z)$ is torsion-free.
@@
We next prove the claims (2) and (3). In the case where $Z=S^{4}$, the claim (2) is trivial. So, we assume $\partial Z$ is non-empty. Since $Z-\mathrm{int}\,D^{4}$ is an invertible cobordism from $S^{3}$ to $\partial Z$, and $H_{2}(Z-\mathrm{int}\,D^{4})$ is isomorphic to $H_{2}(Z)$, the claims (2) and (3) follow from Lemma 6.16. ∎
@@ 8.2. Behavior of genus functions
Now we discuss genus functions. Throughout this subsection, let $X$ be a ((((possibly closed)))) connected 4-manifold, let $L$ be an $S^{2}$-link in $S^{4}$, and let $Z$ be a connected summand or a codimension-00 connected submanifold of the surgered 4-manifold $S^{4}(L)$. We show that genus functions are preserved under connected sums and boundary sums with $Z$.
@@
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the genus functions of $X$ and $X\#Z$. Furthermore, in the case where $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the genus functions.
For the first statement, we identify $H_{2}(X\#Z)$ with $H_{2}(X)\oplus H_{2}(Z)$ by using the isomorphisms $H_{2}(X-\mathrm{int}\,D^{4})\to H_{2}(X)$ and $H_{2}(Z-\mathrm{int}\,D^{4})\to H_{2}(Z)$ induced by the inclusions.
@@
Since $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is diffeomorphic to $Z\#(I\times S^{3})$, Lemma 6.7 shows that $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is a (strongly) quasi-invertible cobordism from $S^{3}$ to the disjoint union $S^{3}\sqcup\partial Z$. The 4-manifold $X\#(Z-\mathrm{int}\,D^{4})$ is constructed from $X-\mathrm{int}\,D^{4}$ by attaching this cobordism, so the inclusion $X-\mathrm{int}\,D^{4}\hookrightarrow X\#(Z-\mathrm{int}\,D^{4})$ preserves the genus functions due to Theorem 7.1. We note that the inclusions $X-\mathrm{int}\,D^{4}\hookrightarrow X$ and $Z-\mathrm{int}\,D^{4}\hookrightarrow Z$ preserve the genus functions by Lemma 7.2. Therefore, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)$ preserves the genus functions.
Now we assume $\partial X$ and $\partial Z$ are non-empty. By Lemma 6.7, any boundary sum $(I\times\partial X)\natural Z$ is a (strongly) quasi-invertible cobordism, where the sum is taken along $1\times\partial X$. Since $X\natural Z$ is constructed from $X$ by attaching this cobordism, the inclusion $X\hookrightarrow X\natural Z$ preserves the genus functions due to Theorem 7.1. ∎
@@
Due to the constructions of $X\#Z$ and $X\natural Z$ in the proof of Theorem 8.4, the claim follows from Proposition 7.6. ∎
Due to the above proposition, torsion-free genus functions are preserved under connected sums and boundary sums with $Z$.
@@
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the torsion-free genus functions of $X$ and $X\#Z$. Furthermore, in the case where both $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the torsion-free genus functions.
Remark 8.7.
The inclusions $H_{2}(Z)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ and $Z\hookrightarrow X\natural Z$ do not necessarily preserve the genus functions. For example, in the case where $X$ is $S^{2}\times S^{2}$ or $S^{2}\times S^{2}-\mathrm{int}\,D^{4}$, we have many counterexamples of $Z$ as easily seen from results of [10], [11]. See also Remark 8.12.
@@ 8.3. Stabilities for algebraic inequivalences of genus functions
Now we discuss stabilities for algebraic inequivalences of genus functions. In the rest of this section, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) connected 4-manifold, let $L_{i}$ be an $S^{2}$-link in $S^{4}$, and let $Z_{i}$ be a connected summand or a codimension-00 connected submanifold of $S^{4}(L_{i})$.
We show that algebraic inequivalences of genus functions are stable under connected sums and boundary sums in the case where $H_{2}(Z_{i})=0$. The theorem below is a restatement of Theorem 1.9.
@@
Under the assumptions, the genus functions of each $X_{i}\#Z_{i}$ and $X_{i}\natural Z_{i}$ are algebraically equivalent to those of $X_{i}$ due to Theorem 8.4. The claims thus immediately follow. ∎
The corresponding results hold for torsion-free genus functions as well.
@@
Due to Corollary 8.6, the claims follow from the same argument as in the proof of Theorem 8.8. ∎
In the case where $b_{2}(Z_{i})\neq 0$, similar stabilities still hold under additional assumptions. The theorem below is a refinement of Theorem 1.10. See Lemma 8.3 for sufficient conditions that each $H_{2}(Z_{i})$ is torsion-free.
@@
Assume that the intersection form of each $X_{i}$ is non-degenerate and that $H_{2}(X_{i})$’s or $H_{2}(Z_{i})$’s are torsion-free. Then the following hold.
$(1)$ Suppose that the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
$(2)$ Suppose that each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty and that the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
@@
Since $H_{2}(S^{4}(L_{i}))=0$ for each $i$, the intersection form of each $Z_{i}$ is the zero-map. Due to Propositions 5.3 and 5.2, the claims thus follow from Theorem 8.4 and Proposition 8.5. ∎
For torsion-free genus functions, the stability holds without the torsion-free assumption.
@@
$(1)$ If the torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ If each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, and the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
@@
Since the intersection form of each $Z_{i}$ is the zero-map, the claims follow from Propositions 5.4 and 8.5 and Corollary 8.6. ∎
@@
Theorems 8.10 and 8.11 do not hold in the case where the intersection forms of $X_{1},X_{2}$ are degenerate by the following trivial reason. Let $X_{1},X_{2}$ be an exotic pair of 4-manifolds with algebraically inequivalent genus functions, and assume they admits embeddings into $S^{4}$. Such examples were produced by Akbulut and the author ([10], [11]), see also Theorem 3.3. Set $Z_{1}=X_{2}$ and $Z_{2}=X_{1}$. In this case $X_{1},X_{2},Z_{1},Z_{2}$ satisfy the assumption of the above theorem except the non-degeneracy assumption, but clearly the conclusions do not hold.
As an application of Theorem 8.8, we show that algebraic equivalent classes of genus functions do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and arbitrarily large $b_{2}$.
Proposition 8.13.
For any closed connected $4$-manifold $X$, there exist infinitely many pairwise non-homeomorphic closed $4$-manifolds with pairwise algebraically equivalent genus functions such that their cohomology rings are isomorphic to those of $X\#S^{1}\times S^{3}$.
@@
Let $K_{1},K_{2},\dots$ be infinitely many $S^{2}$-knots in $S^{4}$ such that their complements $S^{4}-K_{1},S^{4}-K_{2},\dots$ have pairwise non-isomorphic fundamental groups. As is well-known, such families of $S^{2}$-knots exist in abundance and easily produced, for example, as follows. Let $k$ be a knot in $S^{3}$, and let $K$ be the $S^{2}$-knot in $S^{4}$ derived from $k$ by Artin’s spinning construction ([13]). Then $\pi_{1}(S^{4}-K)$ is isomorphic to $\pi_{1}(S^{3}-k)$ (see e.g. [18, Section 2.1]), and thus we obtain a desired infinite family of $S^{2}$-knots by varying a knot $k$. By the Seifert van-Kampen theorem, we see that each $\pi_{1}(S^{4}(K_{i}))$ is isomorphic to $\pi_{1}(S^{4}-K_{i})$. It thus easily follows from Lemma 8.3 that the homology groups of $S^{4}(K_{i})$ are isomorphic to those of $S^{1}\times S^{3}$. Furthermore, we can easily show that the cohomology ring of $S^{4}(K_{i})$ is isomorphic to that of $S^{1}\times S^{3}$ (cf. [17, Section 11]).
Now let $X$ be a closed 4-manifold, and let $Y_{i}$ be the connected sum $X\#S^{4}(K_{i})$. Note that $\pi_{1}(Y_{i})$ is isomorphic to the free product $\pi_{1}(X)*\pi_{1}(S^{4}(K_{i}))$. It thus follows from the Grushko decomposition theorem (see e.g. [60]) that $\pi_{1}(Y_{i})$’s are pairwise non-isomorphic, showing that $Y_{i}$’s are pairwise non-homeomorphic. On the other hand, it is clear that their homology groups, cohomology rings and intersection forms are isomorphic to those of $X\#S^{1}\times S^{3}$. Furthermore, by Theorem 8.4, their genus functions are algebraically equivalent to that of $X$ and hence to that of $X\#S^{1}\times S^{3}$. The claim thus follows. ∎

### v5

@@ 8.1. Submanifolds of $S^{4}(L)$
Before discussing genus functions, we show that various 4-manifolds are realized as codimension-0 submanifolds of the surgered 4-manifold $S^{4}(L)$ (see Definition 6.3) for some $S^{2}$-link $L$ in $S^{4}$. Note that $S^{4}(L)=S^{4}$ in the case where $L$ is an empty set.
@@
Let $L,L^{\prime}$ be $S^{2}$-links in $S^{4}$, and let $V,V^{\prime}$ be connected codimension-0 submanifolds of $S^{4}(L),S^{4}(L^{\prime})$ with non-empty boundaries, respectively. Then, the following hold.
$(1)$ Any boundary sum $V\natural V^{\prime}$ admits an embedding into $S^{4}(L\sqcup L^{\prime})$, where $L\sqcup L^{\prime}$ denotes the split union of $L$ and $L^{\prime}$ in $S^{4}\#S^{4}\cong S^{4}$.
$(2)$ If a $4$-manifold is obtained from $V$ by attaching $1$-handles along a connected component of $\partial V$, then the $4$-manifold admits an embedding into $S^{4}(L)$.
$(3)$ If a $4$-manifold is obtained from $V$ by attaching homotopically canceling pairs of $1$- and $2$-handles, then the $4$-manifold admits an embedding into $S^{4}(L)$.
$(4)$ If a $4$-manifold is obtained from $V$ by attaching $2$-handles along a strongly slice link in $\partial V$ with 0-framings, then the $4$-manifold admits an embedding into $S^{4}(L)$.
$(5)$ Let $\ell$ be a link in $\partial(\natural_{n}S^{2}\times D^{2})$ that bounds a disjoint union of smoothly and properly embedded disks in $\natural_{n}S^{2}\times D^{2}$. Suppose a $4$-manifold is obtained from $\natural_{n}S^{1}\times D^{3}$ by attaching $2$-handles along a link $\ell$ with the framings induced from the disks, where $\ell$ is regarded as a link in $\partial(\natural_{n}S^{1}\times D^{3})$ using an arbitrary diffeomorphism $\partial(\natural_{n}S^{2}\times D^{2})\to\partial(\natural_{n}S^{1}\times D^{3})$. Then the $4$-manifold admits an embedding into $S^{4}$.
@@
Using the above lemma, we can generate numerous examples. Here we provide simple examples of codimension-0 submanifolds of $S^{4}$.
@@
Let $\Sigma_{g}$ denote a closed surface of genus $g$. Then $\Sigma_{g}\times D^{2}$ $(g\geq 0)$, $\natural_{n}S^{1}\times D^{3}$ $(n\geq 0)$, Mazur-type contractible 4-manifolds, their boundary sums, and their connected sums admit embeddings into $S^{4}$ and hence are realized as codimension-0 submanifolds of $S^{4}$. Furthermore, these 4-manifolds with 2-handles attached along strongly slice links with 0-framings are realized as codimension-0 submanifolds of $S^{4}$.
@@
$(2)$ For any codimension-0 submanifold $Z$ of $S^{4}$, $H_{2}(Z)$ is torsion-free.
$(3)$ If the boundary of a codimension-0 submanifold $Z$ of $S^{4}(L)$ is a disjoint union of rational homology $3$-spheres, then $H_{2}(Z)$ is torsion-free.
@@
We next prove the claims (2) and (3). In the case where $Z=S^{4}$, the claim (2) is trivial. So, we assume $\partial Z$ is non-empty. Then, by Lemma 6.7, $Z-\mathrm{int}\,D^{4}$ is a quasi-invertible cobordism from $S^{3}$ to $\partial Z$, and this cobordism is invertible in the case where $Z$ is a submanifold of $S^{4}$. The claims (2) and (3) thus follow from Lemma 6.15, since $H_{2}(Z-\mathrm{int}\,D^{4})$ is isomorphic to $H_{2}(Z)$. ∎
@@ 8.2. Behavior of genus functions
Now we discuss genus functions. Throughout this subsection, let $X$ be a ((possibly closed)) connected 4-manifold, let $L$ be an $S^{2}$-link in $S^{4}$, and let $Z$ be a connected summand or a codimension-0 connected submanifold of the surgered 4-manifold $S^{4}(L)$. We show that genus functions are preserved under connected sums and boundary sums with $Z$.
@@
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the genus functions of $X$ and X#ZX\#Z. Furthermore, in the case where $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the genus functions.
Note that, due to the definition $X\#Z=(X-\mathrm{int}\,D^{4})\cup_{\partial D^{4}}(Z-\mathrm{int}\,D^{4})$, we identify $H_{2}(X\#Z)$ with $H_{2}(X)\oplus H_{2}(Z)$ by using the isomorphisms $H_{2}(X-\mathrm{int}\,D^{4})\to H_{2}(X)$ and $H_{2}(Z-\mathrm{int}\,D^{4})\to H_{2}(Z)$ induced by the inclusions.
@@
Since $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is diffeomorphic to $Z\#(I\times S^{3})$, Lemma 6.7 shows that $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is a strongly quasi-invertible cobordism from $S^{3}$ to the disjoint union $S^{3}\sqcup\partial Z$. The 4-manifold $X\#(Z-\mathrm{int}\,D^{4})$ is constructed from $X-\mathrm{int}\,D^{4}$ by attaching this cobordism, so the inclusion $X-\mathrm{int}\,D^{4}\hookrightarrow X\#(Z-\mathrm{int}\,D^{4})$ preserves the genus functions due to Theorem 7.1. We note that the inclusions $X-\mathrm{int}\,D^{4}\hookrightarrow X$ and $X\#(Z-\mathrm{int}\,D^{4})\hookrightarrow X\#Z$ preserve the genus functions by Lemma 7.2. Therefore, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)$ preserves the genus functions.
Now we assume $\partial X$ and $\partial Z$ are non-empty. By Lemma 6.7, any boundary sum $(I\times\partial X)\natural Z$ is a strongly quasi-invertible cobordism, where the sum is taken along $1\times\partial X$. Since $X\natural Z$ is constructed from $X$ by attaching this cobordism, the inclusion $X\hookrightarrow X\natural Z$ preserves the genus functions due to Theorem 7.1. ∎
@@
Due to the constructions of X#ZX\#Z and $X\natural Z$ in the proof of Theorem 8.4, the claim follows from Theorem 7.7. ∎
Due to the constructions of X#ZX\#Z and $X\natural Z$ in the proof of Theorem 8.4, Corollary 7.8 shows that the torsion-free genus functions are preserved under connected sums and boundary sums with $Z$.
@@
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the torsion-free genus functions of $X$ and X#ZX\#Z. Furthermore, in the case where both $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the torsion-free genus functions.
We remark that the inclusions $H_{2}(Z)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ and $Z\hookrightarrow X\natural Z$ do not necessarily preserve the genus functions. For example, in the case where $X$ is $S^{2}\times S^{2}$ or $S^{2}\times S^{2}-\mathrm{int}\,D^{4}$, we have many counterexamples of $Z$ as easily seen from results of [10], [11]. See also Remark 8.11.
@@ 8.3. Stabilities for algebraic inequivalences of genus functions
Throughout this subsection, for each $i=1,2$, let $X_{i}$ be a ((possibly closed)) connected 4-manifold, let $L_{i}$ be an $S^{2}$-link in $S^{4}$, and let $Z_{i}$ be a connected summand or a codimension-0 connected submanifold of $S^{4}(L_{i})$.
We will show that algebraic inequivalences of genus functions are stable under connected sums and boundary sums. Our results thus show that exotic 4-manifolds with algebraically inequivalent genus functions always yield many more (orientedly) exotic 4-manifolds by taking connected sums and/or boundary sums.
We first show stabilities of algebraic inequivalences in the case where $H_{2}(Z_{i})=0$. Specifically, we prove the following theorem, which is a restatement of Theorem 1.9 and shows the stabilities by taking the contrapositions.
@@
Under the assumptions, Theorem 8.4 shows that the genus functions of each $X_{i}\#Z_{i}$ and $X_{i}\natural Z_{i}$ are algebraically equivalent to that of $X_{i}$, showing the claims. ∎
We also prove stabilities for torsion-free genus functions.
@@
The claims follow from Corollary 8.6 by the same argument as in the proof of Theorem 8.7. ∎
In the case where $b_{2}(Z_{i})\neq 0$, we prove similar stabilities of algebraic inequivalences under additional assumptions. Specifically, we prove the following theorem which is a refinement of Theorem 1.10. See Lemma 8.3 for sufficient conditions that each $H_{2}(Z_{i})$ is torsion-free.
@@
Assume that the intersection form of each $X_{i}$ is non-degenerate and that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are torsion-free. Then the following hold.
$(1)$ Suppose that the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted respectively to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
$(2)$ Suppose that each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty and that the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted respectively to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
@@
Since $H_{2}(S^{4}(L_{i}))=0$ for each $i$, the intersection form of each $Z_{i}$ is the zero-map. Due to Propositions 5.3 and 5.2, the claims thus follow from Theorems 8.4 and 8.5. ∎
For torsion-free genus functions, we can prove stabilities without the torsion-free assumption.
@@
$(1)$ If the torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent, and those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted respectively to $H_{2}(Z_{1})/\mathrm{Tor}$ and $H_{2}(Z_{2})/\mathrm{Tor}$ are algebraically equivalent.
$(2)$ If each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, and the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent, and those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted respectively to $H_{2}(Z_{1})/\mathrm{Tor}$ and $H_{2}(Z_{2})/\mathrm{Tor}$ are algebraically equivalent.
@@
Since the intersection form of each $Z_{i}$ is the zero-map, the claims easily follow from Propositions 5.3 and 5.2, Theorem 8.5 and Corollary 8.6. ∎
@@
Theorems 8.9 and 8.10 do not hold in the case where the intersection forms of $X_{1},X_{2}$ are degenerate by the following simple reason. Let $X_{1},X_{2}$ be an exotic pair of 4-manifolds with algebraically inequivalent genus functions, and assume they admit embeddings into $S^{4}$. Note that there are various examples of such pairs ([10], [11]), see also Theorem 3.7. Set $Z_{1}=X_{2}$ and $Z_{2}=X_{1}$. In this case, $X_{1},X_{2},Z_{1},Z_{2}$ satisfy the assumptions of Theorems 8.9 and 8.10 except the non-degeneracy assumption, but clearly the conclusions do not hold.
As a simple application, we show that algebraic equivalent classes of genus functions do not determine homeomorphism types even within closed 4-manifolds having isomorphic cohomology rings and arbitrarily large $b_{2}$.
Theorem 8.12.
For any closed connected $4$-manifold $X$, there exist infinitely many pairwise non-homeomorphic closed $4$-manifolds with pairwise algebraically equivalent genus functions such that their cohomology rings are isomorphic to that of $X\#S^{1}\times S^{3}$.
@@
Let $K_{1},K_{2},\dots$ be infinitely many $S^{2}$-knots in $S^{4}$ such that their complements $S^{4}-K_{1},S^{4}-K_{2},\dots$ have pairwise non-isomorphic fundamental groups. As is well-known, such families of $S^{2}$-knots exist in abundance and easily produced, for example, as follows. Let $k$ be a knot in $S^{3}$, and let $K$ be the $S^{2}$-knot in $S^{4}$ derived from $k$ by Artin’s spinning construction ([14]). Then $\pi_{1}(S^{4}-K)$ is isomorphic to $\pi_{1}(S^{3}-k)$ (see e.g. [19, Section 2.1]), and thus we obtain a desired infinite family of $S^{2}$-knots by varying a knot $k$. By the Seifert van-Kampen theorem, we see that each $\pi_{1}(S^{4}(K_{i}))$ is isomorphic to $\pi_{1}(S^{4}-K_{i})$. It thus easily follows from Lemma 8.3 that the homology groups of $S^{4}(K_{i})$ are isomorphic to those of $S^{1}\times S^{3}$. Furthermore, we can easily show that the cohomology ring of $S^{4}(K_{i})$ is isomorphic to that of $S^{1}\times S^{3}$ (cf. [18, Section 11]). Note that every $S^{2}$-knot in $S^{4}$ admits a Seifert hypersurface (see e.g. [19, Section 2.5]).
Now let $X$ be a closed connected 4-manifold, and let $Y_{i}$ be the connected sum $X\#S^{4}(K_{i})$. Since each $\pi_{1}(Y_{i})$ is isomorphic to the free product $\pi_{1}(X)*\pi_{1}(S^{4}(K_{i}))$, it follows from the Grushko decomposition theorem (see e.g. [62]) that the fundamental groups $\pi_{1}(Y_{i})$ are pairwise non-isomorphic, showing that the 4-manifolds $Y_{i}$ are pairwise non-homeomorphic. On the other hand, it is clear that their cohomology rings are isomorphic to that of $X\#S^{1}\times S^{3}$. Furthermore, by Theorem 8.4, their genus functions are algebraically equivalent to that of $X$ and hence to that of $X\#S^{1}\times S^{3}$. The claim thus follows. ∎